OpenStudy (anonymous):

f(x)=sen(x+1/2x-3) how to solve this derived...?

OpenStudy (lgbasallote):

\[f(x) = \sin (\frac{x+1}{2x-3})\] correct?

OpenStudy (lgbasallote):

i mean is that the problem?

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To }\mathbb{Open}\textbf{Study}}\] Welcome !!! hope u get answer !!

OpenStudy (angela210793):

what's sen O.o

OpenStudy (anonymous):

spanish version of sin?

mathslover (mathslover):

I think she meant : sin ..

OpenStudy (lgbasallote):

i think it's australian

OpenStudy (angela210793):